Direct labor or machine hours may not be the appropriate cost driver for overhead in all areas of manufacturing due to the complexities of many manufacturing processes. Many companies use activity-based costing (ABC) which uses multiple drivers (items that consume resources) rather than just one driver to apply overhead to their activities. With ABC, a company can use a cost driver that has a direct cause/effect relationship in its applied overhead costs.

Waterways looked into ABC as a method of costing because of the variety of items it produces and the many different activities in which it is involved. The activities listed below area sample of possible cost pools for Waterways.

Assembling Payroll
Billing Plant supervision
Digging trenches Product design
Janitorial Purchasing materials
Machine maintenance Selling
Machine setups Testing
Molding Welding
Packaging

Using the following information, determine the overhead rates and the actual cost assigned for each of the activity cost pools in a possible ABC system for Waterways. (Round answers to 2 decimal places, e.g. 12.25.)

WATERWAYS CORPORATION
Activity Cost
Pools
Cost Drivers Estimated
Overhead
Expected
Use of Cost
Drivers per
Activity
Actual Use of
Drivers
Irrigation installation Labor cost $2,031,480 12,540 12,519
Machining (all machine use) Machine hours 1,816,800 36,336,000 36,337,000
Customer orders Number of orders 32,500 2,500 2,464
Shipping none (direct) N/A N/A traced directly
Design Cost per design 768 8 7
Selling Number of sales calls 312,000 20,800 20,980
WATERWAYS CORPORATION
Activity Cost
Pools
Activity-
Based
Overhead
Rates
Actual
Cost
Assigned
Irrigation installation

$

$

Machining (all machine use)
Customer orders
Design
Selling
Answer #1

Solution :

Calculation of Activity Based Overhead Rate and Allocation of Overhead based on actual activities :

Activity Cost Pool Cost Driver Estimated Overhead Expected Cost Driver Activity Based Over head Rate Actual Cost Driver Actual Cost Assigned
(A) (B) (C) = (A) / (B) (D) (E) = (C) * (D)
Irrigation Installation Labour Cost $ 2,031,480 12,540 $ 162 12,519 $ 2,028,078
Machining Machine Hours $ 1,816,800 36,336,000 $ 0.05 36,337,000 $ 1,816,850
Customer Order No. of Order $ 32,500 2,500 $ 13 2,464 $ 32,032
Shipping None (Direct) N/A N/A NA Traced Directly NA
Design Cost Per Design $ 768 8 $ 96 7 $ 672
Selling No. of Sales Call $ 312,000 20,800 $ 15 20,980 $ 314,700